Reva left this morning, so now just Josh and I from ISF.
Again partly cloudy and not too hot, a little windier today (still out of the south).
Out to the array today between around 11 and 12:
Then back to the base trailer to work on the mice situation. We wiped down all the countertops with bleach and vacuumed the trailer, so now at least we'll know if there are new signs of mice. We also tried adding some steel wool to a couple of the places we think they might be getting in (cable ports, conduit leaving the electrical panel). While vacuuming we did discover one male black widow in the base trailer (behind the bin of spare dsms in the SW corner of the trailer), so keep in mind when moving boxes around. We've also seen them in the sounding/ISFS storage container and in the portapotty. Watch out!!